OpenStudy (anonymous):

lim x to 36.√ x-6/x-36

OpenStudy (anonymous):

is that /x-36 under the root also?

OpenStudy (anonymous):

nah

OpenStudy (anonymous):

oh wait, is it x-6 under the root or just the x?

OpenStudy (anonymous):

the x-6

OpenStudy (anonymous):

man i'm stumped

OpenStudy (zepp):

\[\large \lim_{x\rightarrow 36} \frac{\sqrt{x-6}}{x-36}\]

OpenStudy (anonymous):

yeah deep like that i believe. .

OpenStudy (zepp):

I guess the first step is to rationalize it... @UnkleRhaukus

OpenStudy (zepp):

Wolfram says 36- is -\(\infty\) and 36+ is \(\infty\)

OpenStudy (zepp):

Therefore limit doesn't exist